26 Analysis by a UCGF team has provided the following results 27 28.1 a risk free rate 29 risk aversion coefficient 30market portfolio 6.0% 2.5 18.5% Standard deviation of 31 Calculate the equilibrium risk 2premium Calculate the expected rate of return on the market 35 36 37b another team has determined that the risk aversion coeff is 3.25 what does that suggest for the expected rate return on the market? 40 41describe your conclusion in words 4 c the risk premium of the market has been estimated at 9% 451 the nsk free rate is 3.25% Scenario 1: FG stocks beta is estimated at 145 47 what is the expected rate of return on FG stock? 48 49 50 Scenario 2: the beta was incorrectly computed, it was re-estimated at .95 52 what is the expected rate of return on FG stock in Scenario 2? 51 54 what can be said about FG stock in Scenario 2 vs. Scenario 12 55PLEASE DESCRIBE IN DETAIL sel lil JS Inc. has an expected return of 13% and Beta 12 ss.. . DSCorp has an expected return of 11.55% and Beta of .9 60.1 the markets expected return is 12% and -4% 6ta According to CAPM, which stock is a better buy? 62b What is the alpha of each stock?
Answer #1

II) a) Calculation of Risk Premium and Expected rate of Return :

Risk Premium= Standard deviation of Market Portfolio - Risk Free return

=18.5-6.5=12%

Expected Rate of Return= Rf + Beta of market Portfolio(Rm-Rf)

=6.5 + 2.5(18.5-6.5)= 36.5%

II) b) Calculation of Expected Rate of Return:

Expected Rate of Return= Rf + Beta of Market Portfolio (Rm-Rf)

=6.5 +3.25(18.5-6.5) =45.5%

Conclusion : In the Second Question, the team is expecting more risk on the security because of this Expected rate of Return is also high in this case. If the Investor wants high return on its investment then he have to take more risk.

"High Return brings high risk to the table"

II) c) Scenario I: Risk Free Rate(Rf)=3.25%, Risk Premium= 9%, Beta of Market Portfolio= 1.45

Expected Rate of Return = Rf + Beta(Risk Premium)

=3.25+ 1.45(9)= 16.3%

Scenario II :  Risk Free Rate(Rf)=3.25%, Risk Premium= 9%, Beta of Market Portfolio= .95

Expected Rate of Return = Rf + Beta(Risk Premium)

=3.25+0.95(9)= 11.8%

Conclusion: In Scenario I, the stock is giving high return as beta of security is high but it is wrongly computed and now the beta is less therefore the return is high in this case.

III) A) As Per CAPM,

In case of JS Inc.,

Expected Rate of Return= Rf + Beta ( Rm -Rf)

=4 + 1.2 (12-4) = 13.60%

In Case of DS Corp.,

Expected Rate of Return= Rf +Beta (Rm-Rf)

=4+ .9(12-4) =11.2%

Conclusion : We should recommend to buy JS Inc. as Expected Rate of Return (13.6) is more than Actual Return (13%)

but for DS Corp. Expected Rate of return (11.2%) is less than Actual Return.

Note: Expected Return which is given in question is Actual Return.